Find a number such that 1 less than the number divided by the reciprocal of the number is equal to 1. (Leave the answeres in exact form, that is with the square root.)
Answer by edjones(8007)

let x be the number.
(x-1) / 1/x =1
(x-1) *x=1
x^2-x=1
x^2-x-1=0
x^2-x =1
x^2-x+ 1/4=1+ 1/4
(x- 1/2)^2=5/4
x- 1/2=+-sqrt(5/4)
x=1/2+-sqrt(5)/2
x=(1+sqrt(5))/2, x=(1-sqrt(5))/2
.
Ed